# On the quantum chromatic number of Hamming and generalized Hadamard graphs

### v1

@@ 3 Quantum chromatic number of some Hamming graphs
The goal of this section is to present the proofs of Theorems 1.1 and 1.3.
@@ 3.1 A linear programming approach to orthogonal representations
The following lemma presents an explicit construction of a modulus-one orthogonal representation of Hamming graphs.
@@
The quantity $\xi’(H(n,q,d))$ is upper bounded by the value of any feasible solution to the following linear program:
[…]
@@
If $\frac{(q-1)n}{q}-\frac{\sqrt{(q-1)n}}{q}<d<\frac{(q-1)n}{q}$, then $\xi^{\prime}(H(n,q,d))\leq 2(q-1)^{2}\binom{n}{2}$.
@@
$K_{2}(d)=\frac{q^{2}}{2}d^{2}-\left(q(q-1)n-\frac{q(q-2)}{2}\right)d+\frac{(q-1)^{2}n(n-1)}{2},$
which is negative if and only if $\frac{(q-1)n}{q}-\sqrt{\frac{(q-1)n}{q^{2}}+\frac{(q-2)^{2}}{4}}-\frac{q-2}{2q}<d<\frac{(q-1)n}{q}+\sqrt{\frac{(q-1)n}{q^{2}}+\frac{(q-2)^{2}}{4}}-\frac{q-2}{2q}$. Hence, when $\frac{(q-1)n}{q}-\frac{\sqrt{(q-1)n}}{q}<d<\frac{(q-1)n}{q}$, we have $K_{2}(d)<0$. Therefore, setting $c_{0}=-K_{2}(d)$, $c_{2}=1$, and $c_{i}=0$ for all other $i$ yields a feasible solution.
@@
$\xi^{\prime}(H(n,q,d))\leq-K_{2}(d)+(q-1)^{2}\binom{n}{2}\leq 2(q-1)^{2}\binom{n}{2}$,
where the last inequality holds since $(q-1)^{2}\binom{n}{2}$ is the maximum eigenvalue in absolute value of $H(n,q,2)$ and $K_{2}(d)$ is another eigenvalue of this graph. ∎
@@
where the second inequality holds since $(q-1)^{\lceil z_{1}^{(d)}\rceil}\binom{n}{\lceil z_{1}^{(d)}\rceil}$ is the maximum eigenvalue in absolute value of $H(n,q,\lceil z_{1}^{(d)}\rceil)$ while $K_{\lceil z_{1}^{(d)}\rceil}(d)$ is another eigenvalue of this graph; the last inequality follows from the well-known entropy estimate for Hamming spheres, namely, $(q-1)^{t}\binom{n}{t}\leq q^{h_{q}(\frac{t}{n})n}$. ∎
@@
Theorem 1.1 follows directly from Lemma 2.1 and the combination of Corollaries 3.2, 3.3 and 3.4. ∎
@@ 3.2 Minimum eigenvalue of certain Hamming graphs
The minimum eigenvalue is an important quantity relevant to many combinatorial problems, such as the maximum cut and intersecting families. However, determining it is generally a challenging task. Motivated by semidefinite programming approaches to the max-cut problem on Hamming graphs, Van Dam and Sotirov [14] conjectured that for $d\geq\frac{(q-1)n+1}{q}$, with $d$ even when $q=2$, the minimum eigenvalue of $H(n,q,d)$ is $K_{d}(1)$. Alon and Sudakov [1] proved this for $q=2$ with $n$ large and $d$ fixed. Dumer and Kapralova [7] proved it for $q=2$ and all $n$. Finally, this conjecture was proved by Brouwer et al. [3]. We formulate it as follows:
Lemma 3.5 ([3, Theorem 1.4]).
For $d\geq\frac{(q-1)n+1}{q}$ with $q\geq 3$, the minimum eigenvalue of $H(n,q,d)$ is $K_{d}(1)=-(q-1)^{d}\binom{n}{d}\frac{qd-(q-1)n}{(q-1)n}$.
In the remaining part of this subsection, we focus on the minimum eigenvalue of $H(n,q,d)$ for $d=\frac{(q-1)n}{q}$ with $q\geq 3$. First, we have the following observation.
@@
Let $A$ be the adjacency matrix of the $r$-regular graph $G$. Since each diagonal entry of $A^{2}$ equals $r$, we have $\mathrm{tr}(A^{2})=|V(G)|r$. On the other hand, as the eigenvalues of $A^{2}$ are $\lambda_{i}^{2}$ with multiplicities $m_{i}$, we have $\mathrm{tr}(A^{2})=\sum_{i=1}^{k}m_{i}\lambda_{i}^{2}$. Equating the two expressions gives $\sum_{i}m_{i}\lambda_{i}^{2}=|V(G)|r$. Hence $m_{i}\lambda_{i}^{2}\leq|V(G)|r$ for each $i$, and dividing both sides by $m_{i}r^{2}$ yields $\left(\frac{\lambda_{i}}{r}\right)^{2}\leq\frac{|V(G)|}{rm_{i}}$, as claimed. ∎
@@
For $d=\frac{(q-1)n}{q}$ with $q\geq 3$, the minimum eigenvalue of Hamming graph $H(n,q,d)$ is $K_{d}(2)=-(q-1)^{d}\binom{n}{d}\frac{1}{(q-1)(n-1)}$.
@@
We only consider the following cases: (i) $q\geq 5$ and $n\geq 5$, (ii) $q=4$ and $n\geq 8$, and (iii) $q=3$ and $n\geq 18$, since the remaining finitely many cases can be checked directly.
As discussed in Section 2, the eigenvalues of $H(n,q,d)$ are given by $K_{d}(i)$ for $i=0,1,\dots,n$, with (formal) multiplicities $(q-1)^{i}\binom{n}{i}$. By Lemma 3.6, we have
@@
To show that $K_{d}(2)$ is indeed the minimum eigenvalue, we observe that $K_{d}(0)$ attains the maximum eigenvalue while $K_{d}(1)=0$. Therefore, it suffices to show that for each $3\leq i\leq n$,
@@
Since $K_{d}(2)=-(q-1)^{d}\binom{n}{d}\frac{1}{(q-1)(n-1)}$, after rearranging the inequality it suffices to show that
$(q-1)^{i}\binom{n}{i}\geq\frac{q^{n}(q-1)^{2}(n-1)^{2}}{(q-1)^{d}\binom{n}{d}}$.
The quantity $(q-1)^{i}\binom{n}{i}$ is unimodal in $i$ and attains its maximum at $i=\frac{(q-1)n}{q}$. Therefore,
$(q-1)^{i}\binom{n}{i}\geq\min\!\left\{(q-1)^{3}\binom{n}{3},\,(q-1)^{n}\right\}\geq(q-1)^{3}\binom{n}{3}$,
where the last inequality holds for all three cases (i)–(iii).
Let
$R(q,n)=\frac{(q-1)^{3}\binom{n}{3}(q-1)^{d}\binom{n}{d}}{q^{n}(q-1)^{2}(n-1)^{2}}.$
Using Stirling’s approximation
$\sqrt{2\pi n}\frac{n^{n}}{e^{n}}\leq n!\leq\sqrt{2\pi n}\frac{n^{n}}{e^{n}}e^{\frac{1}{12n}},$
we obtain
$\begin{split}R(q,n)&\geq\frac{(n-2)(q-1)}{6q^{n}}(q-1)^{d}\binom{n}{d}\\ &\geq\frac{(n-2)(q-1)}{6q^{n}}\cdot q^{n}\frac{q}{\sqrt{2\pi n(q-1)}}e^{-\frac{q^{2}}{12(q-1)n}}\\ &\gtrsim_{q}\sqrt{n}.\end{split}$
Hence $R(q,n)$ grows at least on the order of $\sqrt{n}$ and is monotone increasing for large $n$. Consequently, for sufficiently large $n$, we have $R(q,n)\geq 1$. A direct computation further confirms that this already holds at the boundary values: $R(5,5)\geq 1.023\geq 1$, $R(4,8)\geq 1.067\geq 1$, and $R(3,18)\geq 1.108\geq 1$. This completes the proof. ∎
@@
Together with Lemmas 2.2, 3.5 and 3.7, this yields the proof. ∎

### v2

@@ 3 Quantum chromatic number of some Hamming graphs
The goal of this section is to present the proofs of Theorems 1.3 and 1.7.
@@ 3.1 A linear programming approach to orthogonal representations
The following lemma presents an explicit construction of a modulus-one orthogonal representation of Hamming graphs. For the reader’s convenience, we restate it here from the introduction.
@@
The quantity $\xi^{\prime}(H(n,q,d))$ is upper bounded by the objective value of any feasible solution to the following integer linear program:
$\begin{array}[]{lcl}\text{minimize}\quad&\sum\limits_{i=0}^{n}(q-1)^{i}\binom{n}{i}c_{i}&\\[6.0pt] \text{subject to}\quad&\begin{cases}\sum\limits_{i=0}^{n}c_{i}>0,\\ \sum\limits_{i=0}^{n}c_{i}K_{i}(d)=0,\\ c_{0},c_{1},\dots,c_{n}\in\mathbb{Z}_{\geq 0},\end{cases}\end{array}$
where $K_{i}(j)=[z^{i}](1-z)^{j}(1+(q-1)z)^{n-j}$ is the $q$-ary Krawtchouk polynomial. In fact, any feasible solution yields a modulus-one orthogonal representation.
@@
If $\frac{(q-1)n}{q}>d\geq\frac{(q-1)n-(\frac{q-2}{2})-\sqrt{(q-1)n+(\frac{q-2}{2})^{2}}}{q}$, then
$\xi^{\prime}(H(n,q,d))\leq\left((q-1)n-\frac{q-2}{2}\right)qd-\frac{1}{2}q^{2}d^{2}.$
@@
$K_{2}(d)=\frac{1}{2}q^{2}d^{2}-\left((q-1)n-\frac{q-2}{2}\right)qd+(q-1)^{2}\binom{n}{2}$,
which is negative if and only if $\frac{(q-1)n-(\frac{q-2}{2})-\sqrt{(q-1)n+(\frac{q-2}{2})^{2}}}{q}<d<\frac{(q-1)n-(\frac{q-2}{2})+\sqrt{(q-1)n+(\frac{q-2}{2})^{2}}}{q}$. Hence, when $\frac{(q-1)n}{q}>d\geq\frac{(q-1)n-(\frac{q-2}{2})-\sqrt{(q-1)n+(\frac{q-2}{2})^{2}}}{q}$, we have $K_{2}(d)\leq 0$. Therefore, setting $c_{0}=-K_{2}(d)$, $c_{2}=1$, and $c_{i}=0$ for all other $i$ yields a feasible solution.
@@
$\xi^{\prime}(H(n,q,d))\leq-K_{2}(d)+(q-1)^{2}\binom{n}{2}=\left((q-1)n-\frac{q-2}{2}\right)qd-\frac{1}{2}q^{2}d^{2}$
@@
where the second inequality holds since $(q-1)^{\lceil z_{1}^{(d)}\rceil}\binom{n}{\lceil z_{1}^{(d)}\rceil}$ is the maximum eigenvalue in absolute value of $H(n,q,\lceil z_{1}^{(d)}\rceil)$ while $K_{\lceil z_{1}^{(d)}\rceil}(d)$ is another eigenvalue of this graph; the last inequality follows from the well-known entropy estimate for Hamming spheres, namely, $(q-1)^{t}\binom{n}{t}\leq q^{H_{q}(\frac{t}{n})n}$, where the estimate holds whenever $t\leq\frac{(q-1)n}{q}$. ∎
@@
Theorem 1.3 follows directly from Lemma 1.1 and the combination of Corollaries 3.1, 3.2 and 3.3. ∎
@@ 3.2 Minimum eigenvalue of certain Hamming graphs
For $d\geq\frac{(q-1)n+1}{q}$, Brouwer et al. [6] showed that:
Lemma 3.4 ([6, Theorem 1.4]).
For $d\geq\frac{(q-1)n+1}{q}$ with $d$ even when $q=2$, the minimum eigenvalue of $H(n,q,d)$ is $K_{d}(1)$.
In the remaining part of this subsection, we focus on the minimum eigenvalue of $H(n,q,d)$ in the regime where $d$ lies slightly below the threshold $\frac{(q-1)n}{q}$, with $q$ fixed and $n$ sufficiently large. First, we have the following observation.
@@
Let $A$ be the adjacency matrix of the $r$-regular graph $G$. Since each diagonal entry of $A^{2}$ equals $r$, we have $\mathrm{tr}(A^{2})=|V(G)|r$. On the other hand, as the eigenvalues of $A^{2}$ are $\lambda_{i}^{2}$ with multiplicities $m_{i}$, we have $\mathrm{tr}(A^{2})=\sum_{i=1}^{k}m_{i}\lambda_{i}^{2}$. Equating the two expressions gives $\sum_{i}m_{i}\lambda_{i}^{2}=|V(G)|r$. Hence $m_{i}\lambda_{i}^{2}\leq|V(G)|r$ for each $i$, and dividing both sides by $m_{i}r^{2}$ yields $(\frac{\lambda_{i}}{r})^{2}\leq\frac{|V(G)|}{rm_{i}}$, as claimed. ∎
For the reader’s convenience, here we restate the following lemma from the introduction.
@@
For $\frac{(q-1)n}{q}\geq d\geq\frac{(q-1)n-\frac{q-2}{2}-\sqrt{(q-1)n+(\frac{q-2}{2})^{2}-(q-1)}+1}{q}$, with $d$ even when $q=2$ and $n$ sufficiently large, the minimum eigenvalue of $H(n,q,d)$ is $K_{d}(2)$.
@@
As discussed in Section 2, the eigenvalues of $H(n,q,d)$ are given by $K_{d}(i)$ for $i=0,1,\dots,n$, with (formal) multiplicities $(q-1)^{i}\binom{n}{i}$.
We begin by showing that $K_{d}(2)\leq K_{d}(i)$ for $i\leq 3$. By reciprocity, $K_{d}(i)=(q-1)^{d}\binom{n}{d}\frac{K_{i}(d)}{(q-1)^{i}\binom{n}{i}}$, and therefore the inequality $K_{d}(i)\leq K_{d}(j)$ is equivalent to $\frac{K_{i}(d)}{(q-1)^{i}\binom{n}{i}}\leq\frac{K_{j}(d)}{(q-1)^{j}\binom{n}{j}}$.
Observe that
[…]
Clearly, $K_{0}(d)>0$ for all $d$, $K_{1}(d)\geq 0$ for $d\leq\frac{(q-1)n}{q}$, and $K_{2}(d)$ is negative if and only if $\frac{(q-1)n-\frac{q-2}{2}-\sqrt{(q-1)n+(\frac{q-2}{2})^{2}}}{q}<d<\frac{(q-1)n-\frac{q-2}{2}+\sqrt{(q-1)n+(\frac{q-2}{2})^{2}}}{q}$. Therefore, under the assumptions of the lemma on $d$, we have $K_{2}(d)<0$.
Next, we compare $\frac{K_{2}(d)}{(q-1)^{2}\binom{n}{2}}$ and $\frac{K_{3}(d)}{(q-1)^{3}\binom{n}{3}}$. Define
$f(z)=\frac{K_{2}(z)}{(q-1)^{2}\binom{n}{2}}-\frac{K_{3}(z)}{(q-1)^{3}\binom{n}{3}}$.
Observe that f(0)=11=0f(0)=1-1=0, and hence the monomial $z$ divides $f(z)$. Consequently, $f(z)/z$ is a quadratic polynomial, whose roots are easily computed to be $\frac{(q-1)n-\frac{q-2}{2}\pm\sqrt{(q-1)n+(\frac{q-2}{2})^{2}-(q-2)}+1}{q}$. Since $f(z)$ is a cubic polynomial with positive leading coefficient, it follows that under the assumptions of the lemma on $d$, we have $f(d)\leq 0$. Thus, $K_{d}(2)\leq K_{d}(i)$ for $i\leq 3$.
It remains to show that $K_{d}(2)\leq K_{d}(i)$ for $4\leq i\leq n$. Note that when $q=2$ and $d$ is even, we have $K_{d}(i)=K_{d}(n-i)$. Therefore, in this case, it suffices to consider $4\leq i\leq\frac{n}{2}$.
In fact, we prove the following stronger statement:
  • 1.
    for $q=2$ and $4\leq i\leq\frac{n}{2}$;
  • 2.
    for $q\geq 3$ and $4\leq i\leq n$,
we have $|K_{d}(2)|\geq|K_{d}(i)|$ for sufficiently large $n$.
By Lemma 3.5, we have
@@
Thus, it suffices to show that
@@
On the one hand, we derive an upper bound for the left-hand side. Since $(q-1)^{i}\binom{n}{i}$ is unimodal in $i$ and attains its maximum at $i=\frac{(q-1)n}{q}$, we conclude that for sufficiently large $n$, $(q-1)^{i}\binom{n}{i}\geq(q-1)^{4}\binom{n}{4}$ for $q=2$ and $4\leq i\leq\frac{n}{2}$, and for $q\geq 3$ and $4\leq i\leq n$. Moreover, by a standard local limit estimate for the binomial distribution, for $d=\frac{(q-1)n}{q}-O(\sqrt{n})$, we have $(q-1)^{d}\binom{n}{d}\sim\frac{q^{n}}{\sqrt{n}}$. Therefore, under the assumptions of the lemma on $d$, we have
$\frac{q^{n}}{(q-1)^{d}\binom{n}{d}(q-1)^{i}\binom{n}{i}}\lesssim_{q}\frac{q^{n}}{\frac{q^{n}}{\sqrt{n}}(q-1)^{4}\binom{n}{4}}\lesssim_{q}n^{-3.5}.$
On the other hand, under the assumptions of the lemma on $d$, the quadratic polynomial $K_{2}(d)$ attains its maximum at $d=\frac{(q-1)n-\frac{q-2}{2}-\sqrt{(q-1)n+(\frac{q-2}{2})^{2}-(q-1)}+1}{q}$, with value $\frac{q-2}{2}-\sqrt{(q-1)n+(\frac{q-2}{2})^{2}-(q-1)}$. Thus,
$\left(\frac{K_{d}(2)}{(q-1)^{d}\binom{n}{d}}\right)^{2}=\left(\frac{K_{2}(d)}{(q-1)^{2}\binom{n}{2}}\right)^{2}\geq\left(\frac{-\frac{q-2}{2}-\sqrt{(q-1)n+(\frac{q-2}{2})^{2}-(q-1)}}{(q-1)^{2}\binom{n}{2}}\right)^{2}\gtrsim_{q}n^{-3}.$
Hence, we conclude that $K_{d}(2)\leq K_{d}(i)$ for all $i\geq 4$, which completes the proof. ∎
@@
Together with Lemmas 1.2, 3.4 and 1.6, this yields the proof. ∎